y=log(3)(x) for 1≤x≤243 graph it
Answer by lwsshak3(11628)

y=log(3)(x) for 1≤x≤243 graph it
***
f(x)=log3(x)
y=log3(1)=0 (3^0=1)
y=log3(243)=5 (3^5=243)
y=log3(81)=4 (3^4=81)
y=log3(27)=3 (3^3=27
..
You now have 4 points with which you can draw the graph: (1,0), (27,3), (81,4) and (243,5)